You can buy a machine for $100,000 that will produce a net income, after operating expenses, of $10,000 per year. If you plan to keep the machine for four years, what must the market (resale) value be at the end of four years to justify the investment? Assume interest rate at 15%.
1
Expert's answer
2020-05-13T11:07:40-0400

Alternative investments=x

"x=100,000*1.15^4=174,900.625"

A net income for four-year=y

"y=10,000*(1+0.15)^3+10,000*(1+0.15)^2+10,000*(1+0.15)+10,000=49,933.75"

The market (resale) value should be at the end of four-year="174,900.625-49,933.75=124,966.875"